# Thermal avalanches in isolated many-body localized systems

### v2

@@ Abstract
Introduction.– The presence of disorder in interacting many-body systems results in the emergence of an interesting phase of matter known as many-body localization (MBL) which affects the entire spectrum of the system [1, 2, 3, 4, 5, 6]. Over the past decade, many distinctive features of MBL have been identified, including the breakdown of eigenstate thermalization hypothesis [7, 8, 9], the emergence of quasi-local integrals of motion [10, 11], Poisson-like level statistics [12, 13, 14, 15, 16], the emergence of mobility edge [15, 17, 18], area-law entangled eigenstates [19, 20], and logarithmically slow entanglement growth [21, 22, 23, 24, 25, 26, 27]. Thanks to recent advancement in quantum technologies, various aspects of MBL have also been tested experimentally on various physical platforms [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. Nonetheless, the stability of the MBL phase in the presence of a small thermal region in the system has been challenged by the avalanche theory [52, 53]. Thus, a key open question is to identify the criteria, likely dependent on both the size of the thermal region and the disorder strength, that trigger an avalanche, thereby destabilizing the MBL phase.
@@
Analytical studies and renormalization group approaches indicate that quantum avalanches can arise when small ergodic regions or thermal inclusions act as local baths, gradually absorbing the surrounding localized degrees of freedom [52, 54, 55, 56, 57, 52, 53, 53, 58, 59]. In this scenario, MBL is not a genuine thermodynamic phase but only a prethermal regime that breaks down at sufficiently long times [60, 61, 62, 63] or large system sizes [64]. The avalanche mechanism has been investigated via two primary frameworks: (i) isolated systems, where weak (or zero) disorder defines the thermal region [65, 64, 57]; and (ii) open quantum systems, where the thermal region couples to an external bath [66, 67, 68, 69]. Experimental evidence for avalanche-like behavior has recently been observed in cold-atom systems [70].
@@
Figure 1: (a) Schematic of our model system. While the dark points are exposed to strong random disorder, the light points are subjected to a weak random disorder of strength $w{=}0.5$. (b) Graphical visualization of the phase diagram as a function of disorder strength $W$ and the size of the weak-disorder region $P{/}L$. The solid (dashed) line indicates the phase boundary controlled by the critical $W$ ($P{/}L$). The intermediate area gives a rough approximation of the region wherein the MBL is unstable in the thermodynamic limit.
@@
In contrast, a parallel line of research argues that MBL can persist even in systems with avalanche-inducing features, provided that certain conditions are satisfied [59, 65, 67, 66, 71, 72]. These studies show that localization instability is not universal and may depend sensitively on system architecture [56, 73, 71, 72, 74], interaction range [75, 76], or even the characteristics of the thermal regions [77]. For instance, while a fixed-size bath that is described by a suitable random matrix predicts the instability of the MBL [58], the presence of the same size thermal region in the Heisenberg chain may not destabilize the MBL phase [65]. Indeed, to resolve these conflicts a systematic analysis of MBL instability under the presence of a thermal region is highly desirable. Here, we address this issue by numerically investigating the stability of the MBL in an isolated Heisenberg spin chain of size $L$ subjected to a random magnetic field. We intentionally embed a thermal region of $P$ spins with weak-disorder and analyze the system’s static and dynamical behavior using entanglement entropy and the gap ratio. Our results show that avalanches occur in strongly disordered systems only when the thermal region scales with system size and particularly exceeds a threshold value of $P{/}L{>}0.1$. As $P$ increases, an intermediate phase emerges between the MBL and ergodic regimes, manifesting in both static and dynamic properties of the system. However, this phase is expected to vanish in the thermodynamic limit. Fig. 1 (b) presents a graphical visualization of the extracted phase diagram as a function of disorder strength $W$ and the relative size of the thermal region $P{/}L$. The solid (dashed) line approximates the phase boundaries controlled by disorder strength $W$ (relative size of the thermal region $P{/}L$).
@@
Model and methodology.– We consider a Heisenberg spin-$1/2$ chain of size $L$ in the presence of disordered magnetic fields as
@@
where, $J{=}1$ is the exchange coupling, $\boldsymbol{\sigma}_{i}{=}(\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z})$ is the vector of Pauli matrices, $r_{i}$’s are random numbers drawn from a uniform distribution $[-1,1]$, $w{=}0.5$ is a fixed weak disorder strength which affects the $P$ spins at the center of the chain (which we call it region $\mathcal{P}$) and $W$ is a controllable disorder strength which affects the spins out of the region $\mathcal{P}$. See Fig. 1(a) for the schematic of the system. Since the Hamiltonian conserves total magnetization $\sigma^{z}_{\mathrm{tot}}{=}\sum_{i}\sigma_{i}^{z}$, we restrict our numerical analysis to the special sector of the spectrum which is spanned by a set of eigenstates $\{|E_{k}\rangle\}$ such that $\left\langle E_{k}\right|\sigma^{z}_{\text{total}}\left|E_{k}\right\rangle{=}0$. We investigate the physics of the system by controlling two parameters, namely $W$ and $P$. Without the weak-disorder region $\mathcal{P}$, namely for the fully disordered chain, the system exhibits a phase transition at a specific disorder strength $W{=}W_{c}$ between ergodic and MBL phases. In the following, we investigate the MBL transition and its stability for $P{>}0$ through static and dynamical analysis.
@@
Static analysis.– In the static case, we calculate two standard quantities, namely the gap ratio (GR) and bipartite entanglement entropy (EE), for 50 eigenenergies $E_{k}$ and the corresponding eigenstates ${\left|E_{k}\right\rangle}$ in the mid-spectrum $\varepsilon{=}0.5$ of the Hamiltonian $H$, extracted using the exact diagonalization method after normalizing the energy spectrum as $\varepsilon{=}{(E_{k}-E_{\min})/(E_{\max}-E_{\min}})$ wherein $E_{\max}$ ($E_{\min}$) is the largest (smallest) eigenenergy of the system. For an ordered set of eigenenergies $\{E_{k}\}$, with energy gaps between consecutive levels $\delta_{k}{=}E_{k+1}{-}E_{k}$, the ratio $r_{k}{=}\text{min}(\delta_{k+1},\delta_{k})/\text{max}(\delta_{k+1},\delta_{k})$ and its average $\overline{\langle r\rangle}$ over energy levels and sample realizations can diagnose a global change in spectral statistics. Strong repulsion between neighboring levels in the ergodic phase yields Gaussian distribution with $\overline{\langle r\rangle}{\simeq}0.5307$. However, the MBL phase exhibits Poisson distribution with $\overline{\langle r\rangle}{\simeq}0.3863$. The EE for each eigenstate $\left|E_{k}\right\rangle$ is calculated by dividing the chain into two halves $A$ and $B$ regardless of weak-disorder region $\mathcal{P}$, and using $S(|E_{k}\rangle){=}{-}\text{Tr}[\rho_{A}^{k}\ln\rho_{A}^{k}]$, where $\rho_{A}^{k}{=}\text{Tr}_{B}[{\left|E_{k}\right\rangle\left\langle E_{k}\right|}]$ is the reduced density matrix of the subsystem $A$ obtained after tracing out the subsystem $B$. In the ergodic phase, eigenstates are expected to resemble random pure states, exhibiting volume-law EE consistent with Page entropy $S_{\mathrm{P}}{=}{\sum_{k=\mathcal{H}_{B}+1}^{\mathcal{H}_{A}\mathcal{H}_{B}}\ \frac{1}{k}{-}\frac{(\mathcal{H}_{A}-1)}{2\mathcal{H}_{B}}}$, wherein $\mathcal{H}_{A}$ ($\mathcal{H}_{B}$) is the dimension of the Hilbert space associated with the subsystem $A$ ($B$[78, 79]. Studies indicate that highly excited eigenstates in the MBL phase obey an area law for bipartite entanglement, due to the emergent integrability [80, 81]. Therefore, averaged EE over eigenstates and disorder realizations $\overline{\langle S\rangle}$, results in $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}1$ in the ergodic phase, and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}0$ in the MBL phase.
@@
Figure 2: (a)-(b) the averaged EE and averaged GR, respectively, as a function of disorder strength $W$ in fully disordered chains of different sizes. (c)-(d) $\overline{\langle S\rangle}/S_{\mathrm{P}}$ and $\overline{\langle r\rangle}$, respectively, as a function of $P{/}L$ in systems with varying sizes exposed to strong disorder $W{=}9$. The results for EE and GR are obtained using $1000{-}2000$ and $10000{-}25000$ realizations, respectively.
@@
The interplay between the MBL and ergodic phases is determined by two competing factors: (i) the disorder strength $W$; and (ii) the size of the thermal region $\mathcal{P}$, indicated by $P$ spins. While increasing $W$ tends to localize the system, increasing $P$ weakens the effect of disorder and tends to thermalize the system. In order to investigate these effects, we first fix $P$ and vary $W$. By increasing $W$, the system goes through a phase transition from the ergodic to the localized phase. Figure 2(a) depicts the behavior of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ as a function of the disorder strength $W$, for different system sizes $L{\in}[10,{\cdots},16]$, when $P{=}0$. As the figure shows, the system is ergodic for small disorders and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}1$, indicating that the EE fulfills the volume-law entanglement, while increasing disorder $W$ leads to localization and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}0$ [82]. The transition between these two phases happens at $W{=}W_{c}$. Thanks to scale invariant behavior at the phase transition point, one can utilize finite-size scaling analysis [83, 84, 85] to extract the critical value $W_{c}{\simeq}4.674$ as the prediction for the transition point in the thermodynamic limit, for a system with $P{=}0$. Similar analysis can be performed for the GR which is shown in Fig. 2(b). Decreasing GR from $\overline{\langle r\rangle}{\simeq}0.5307$ to $\overline{\langle r\rangle}{\simeq}0.3863$ clearly demonstrates the transition from the ergodic to the MBL phase. Here, finite-size scaling analysis results in $W_{c}{\simeq}4.882$ (for $P{=}0$), which is close to the value identified by the EE. Repeating the above analysis, using either EE or GR, we extract $W_{c}$ for various values of $P$ and report the results in the Supplementary Materials (SM).
@@
Alternatively, one can fix disorder strength $W$ to a large value, enough for localizing the system for $P{=}0$, and then increase $P$ to gradually thermalize the system. In Fig. 2(c), we plot $\overline{\langle S\rangle}/S_{\mathrm{P}}$ as a function of $P{/}L$, keeping $W{=}9$ fixed, for various system sizes. Several important features are observed. First, regardless of the system size, MBL can survive the presence of a small weak-disorder region $\mathcal{P}$, evidenced by the early plateau of the curves. Second, after the initial plateau when $P{/}L{>}0.1$, the value of EE increases linearly by increasing $P/L$ such that the system eventually thermalizes to a volume law state. Third, the rise of EE for $P{/}L{>}0.1$, is sharpened as the system size increases suggesting that in the thermodynamic limit, thermalization occurs abruptly. In Fig. 2(d), we depict $\overline{\langle r\rangle}$ versus $P{/}L$ in systems with different sizes, keeping the disorder strength fixed at $W{=}9$. The results provide further qualitative confirmation for the findings from the EE. Initially, the GR increases very slowly indicating the resilience of the MBL phase against small weak-disorder regions in all considered system sizes. The growth of the GR increases significantly as $P{/}L$ increases, again indicating the instability of the MBL for reasonably large weak-disorder regions. Similar to the EE case, the growth rate sharpens as the system size increases. The above observations, for both GR and EE, lead us to the following conjecture: to trigger an avalanche and destabilize the MBL phase, the ratio $P/L$ should exceed a critical threshold, i.e. $P/L{\gtrsim}0.1$.
Figure 3: (a) averaged EE $\overline{\langle S\rangle}{/}S_{P}$ and (b) averaged GR $\overline{\langle r\rangle}$ as a function of $W$ and $P{/}L$ in a system of size $L{=}16$. In both panels, the solid circles represent the estimated $W_{c}^{*}(L_{av})$ for fixed $P$’s. The other colorful markers represent the extracted $(P{/}L_{av})_{c}^{*}$ for fixed $W$’s. As graphically visualized in Fig.1(b), three distinct regions are classified as $1$ ergodic, $2$ intermediate, and $3$ localized.
@@
To illustrate a complete picture of the effect of both disorder strength and the weak-disorder region, in Fig. 3(a) we plot $\overline{\langle S\rangle}/S_{\mathrm{P}}$ versus $W$ and $P{/}L$, obtained in a system of size $L{=}16$. The light yellow area represents the ergodic phase and the dark blue determines the MBL spot. As mentioned above, there are two ways to obtain the phase boundary (i) we can fix $P$ and vary $W$; or (ii) we can fix disorder strength $W$ and vary $P$. In the first approach, we find the intersection of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ curves for system of sizes $L$ and $L{+}2$ when they are plotted as a function of $W$ for a given $P$, e.g. see Fig. 2(a) for $P{=}0$. The intersection point $W_{c}^{*}(L_{av})$, with $L_{av}{=}L{+}1$ is taken as the finite-size transition point, at a given $P$, see the SM for more details. The results for $W_{c}^{*}(L_{av})$ are shown in Fig. 3(a) with solid circles. Different lines correspond to different sets of system sizes that are used to extract $W_{c}^{*}(L_{av})$. The smooth dependence of the EE on $L$ and $W$, especially at $P{=}0$, guarantees that $W_{c}^{*}(L_{av})$ is a lower bound for the MBL transition point in the thermodynamic limit, namely $W_{c}$. As it is evident, by increasing $P$, the phase boundary initially remains straight and then bends after $P{/}L{\gtrsim}0.25$. This shows that the MBL phase is indeed stable against the presence of small thermal regions. On the other hand, one can also determine the phase boundary between the ergodic and the MBL phases by fixing $W$ and varying $P$. For every given $W$, we determine the transition point $(P{/}L_{av})_{c}^{*}$ by finding the intersection point of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ curves for systems of size $L$ and $L{+}2$, when they are plotted as a function of $P/L$ for a fixed value of $W$, e.g. see Fig. 2(c) for $W{=}9$. The results are shown by stars in Fig. 3(a). Interestingly, the phase boundary between the MBL and the ergodic phase is found to be different depending on the two above approaches. This suggests that there exists an intermediate region between the ergodic and the MBL phase, which we cannot decisively label as MBL or ergodic. We call this area in the phase diagram the intermediate phase, which is also schematically shown in Fig. 1(b). As discussed for Figs. 2(c) and (d) (see also the SM for other $W$ values), this intermediate phase tends to vanish in the thermodynamic limit as the localization to ergodic transition accelerates by increasing the system size.
@@
Having elucidated the EE behavior for inspecting the stability of MBL, we now focus on the GR. In Fig. 3 (b) we report the obtained $\overline{\langle r\rangle}$ as a function of $W$ and $P{/}L$ for a system of size $L{=}16$. By repeating the above analysis to determine the phase boundaries, one again finds different phase boundaries depending on whether $W$ is varied while $P/L$ is kept fixed or $P/L$ is varied while $W$ is fixed. Similar to the EE, the GR also shows relative stability for the MBL phase for small values of $P/L$, which is evident in the bending of the phase boundary as $P/L$ increases. Clearly, the GR analysis also leaves an intermediate phase between the ergodic and the MBL phases, which is not fully decisive. Therefore, both EE and GR predict an intermediate phase for finite systems which lies between the MBL and the ergodic phases.
Figure 4: (a) The dynamic of EE in a system of size $L{=}16$ for different values of $P$, when the system with $P{=}0$ is determined as localized in the presence of strong disorder $W{=}9$. (b) the saturated values of the EE, namely $\overline{S(t{\rightarrow)\infty}}{/}S_{P}$, as a function of $P{/}L$ when the size of the system varies and $W{=}9$. The results are obtained using $200{-}400$ realizations.

### v3

@@ I Introduction
The presence of disorder in interacting many-body systems results in the emergence of an interesting phase of matter known as many-body localization (MBL) which affects the entire spectrum of the system [1, 2, 3, 4, 5, 6, 7]. Over the past decade, many distinctive features of MBL have been identified, including the breakdown of eigenstate thermalization hypothesis [8, 9, 10], the emergence of quasi-local integrals of motion [11, 12], Poisson-like level statistics [13, 14, 15, 16, 17], the emergence of mobility edge [16, 18, 19], area-law entangled eigenstates [20, 21], and logarithmically slow entanglement growth [22, 23, 24, 25, 26, 27, 28]. Thanks to recent advances in quantum technologies, various aspects of MBL have also been tested experimentally on various physical platforms [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. Nonetheless, the stability of the MBL phase in the presence of a small thermal region in the system has been challenged by the avalanche theory [53, 54]. Thus, a key open question is to identify the criteria, likely dependent on both the size of the thermal region and the disorder strength, that trigger an avalanche, thereby destabilizing the MBL phase.
@@
Analytical studies and renormalization group approaches indicate that quantum avalanches can arise when small ergodic regions or thermal inclusions act as local baths, gradually absorbing the surrounding localized degrees of freedom [53, 55, 56, 57, 58, 53, 54, 54, 59, 60]. In this scenario, MBL is not a genuine thermodynamic phase but only a prethermal regime that breaks down at sufficiently long times [61, 62, 63, 64] or large system sizes [65]. The avalanche mechanism has been investigated via two primary frameworks: (i) isolated systems, where weak (or zero) disorder defines the thermal region [66, 65, 58, 67]; and (ii) open quantum systems, where the thermal region couples to an external bath [68, 69, 70, 71]. Experimental evidence for avalanche-like behavior has recently been observed in cold-atom systems [72].
@@
Figure 1: (a) Schematic of our model system. While the dark points are exposed to strong random disorder, the light points are subjected to a weak random disorder of strength w=0.5w{=}0.5. (b) Graphical visualization of the phase diagram as a function of disorder strength $W$ and the size of the weak-disorder region P/LP{/}L. The solid (dashed) line indicates the phase boundary controlled by the critical $W$ (P/LP{/}L). The intermediate area gives a rough approximation of the region wherein the MBL is unstable in the thermodynamic limit. (c) Schematic of two different types of chain partitioning for calculation of the entanglement entropy.
@@
In contrast, a parallel line of research argues that MBL can persist even in systems with avalanche-inducing features, provided that certain conditions are satisfied [60, 66, 69, 68, 73, 74]. These studies show that localization instability is not universal and may depend sensitively on system architecture [57, 75, 73, 74, 76], interaction range [77, 78], or even the characteristics of the thermal regions [79]. For instance, while a fixed-size bath that is described by a suitable random matrix predicts the instability of the MBL [59], the presence of the same size thermal region in the Heisenberg chain may not destabilize the MBL phase [66]. Indeed, to resolve these conflicts a systematic analysis of MBL instability under the presence of a thermal region is highly desirable.
Here, we address this issue by numerically investigating the stability of the MBL in an isolated Heisenberg spin chain of size $L$ subjected to a random magnetic field. We intentionally embed a thermal region of $P$ spins with weak-disorder and analyze the system’s static and dynamical behavior using entanglement entropy and the gap ratio. Our results show that avalanches occur in strongly disordered systems only when the thermal region scales with system size and particularly exceeds a threshold value of P/L>0.1P{/}L{>}0.1. As $P$ increases, an intermediate phase emerges between the MBL and ergodic regimes, manifesting in both static and dynamic properties of the system. However, this phase is expected to vanish in the thermodynamic limit. Fig. 1(b) presents a graphical visualization of the extracted phase diagram as a function of disorder strength $W$ and the relative size of the thermal region P/LP{/}L. The solid (dashed) line approximates the phase boundaries controlled by disorder strength $W$ (relative size of the thermal region P/LP{/}L).
II Model and methodology
@@
We consider a Heisenberg spin-$1/2$ chain of size $L$ in the presence of disordered magnetic fields as
@@
where, J=1J{=}1 is the exchange coupling, $\boldsymbol{\sigma}_{i}{=}(\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z})$ is the vector of Pauli matrices, $r_{i}$’s are random numbers drawn from a uniform distribution [1,1][-1,1], w=0.5w{=}0.5 is a fixed weak disorder strength which affects the $P$ spins at the center of the chain (which we call it region $\mathcal{P}$) and $W$ is a controllable disorder strength which affects the spins out of the region $\mathcal{P}$. See Fig. 1(a) for the schematic of the system. Since the Hamiltonian conserves total magnetization $\sigma^{z}_{\mathrm{tot}}{=}\sum_{i}\sigma_{i}^{z}$, we restrict our numerical analysis to the special sector of the spectrum which is spanned by a set of eigenstates $\{|E_{k}\rangle\}$ such that $\left\langle E_{k}\right|\sigma^{z}_{\text{tot}}\left|E_{k}\right\rangle{=}0$. We investigate the physics of the system by controlling two parameters, namely $W$ and $P$. Without the weak-disorder region $\mathcal{P}$, namely for the fully disordered chain, the system exhibits a phase transition at a specific disorder strength $W{=}W_{c}$ between ergodic and MBL phases. In the following, we investigate the MBL transition and its stability for $P{\geq}0$ through static and dynamical analysis.
III Static analysis
@@
In the static case, we calculate two standard quantities, namely the gap ratio (GR) and bipartite entanglement entropy (EE), for 50 eigenenergies $E_{k}$ and the corresponding eigenstates $\left|E_{k}\right\rangle$} in the mid-spectrum $\varepsilon{=}0.5$ of the Hamiltonian $H$, extracted using the exact diagonalization method after normalizing the energy spectrum as $\varepsilon{=}{(E_{k}-E_{\min})/(E_{\max}-E_{\min}})$ wherein $E_{\max}$ ($E_{\min}$) is the largest (smallest) eigenenergy of the system. For an ordered set of eigenenergies $\{E_{k}\}$, with energy gaps between consecutive levels $\delta_{k}{=}E_{k+1}{-}E_{k}$, the ratio $r_{k}{=}\text{min}(\delta_{k+1},\delta_{k})/\text{max}(\delta_{k+1},\delta_{k})$ and its average $\overline{\langle r\rangle}$ over energy levels and sample realizations can diagnose a global change in spectral statistics. Strong repulsion between neighboring levels in the ergodic phase yields Gaussian distribution with $\overline{\langle r\rangle}{\simeq}0.5307$. However, the MBL phase exhibits Poisson distribution with $\overline{\langle r\rangle}{\simeq}0.3863$. As the EE behavior depends on the choice of partitioning, for each eigenstate $\left|E_{k}\right\rangle$, the EE can be calculated by partitioning the chain in different ways. In this work, we focus on the EE for two types of partitioning presented in Fig. 1(c). In the first case, we cut the chain from the midpoint in a way that the subsystems $A$ and $B$ are equal in size. In the second scenario, we separate the strongly disordered region on the most left side from the rest of the chain, resulting in non-equal subsystems $A$ and $B$. In both cases, the EE is defined as $S(\left|E_{k}\right\rangle){=}{-}\text{Tr}[\rho_{A}^{k}\ln\rho_{A}^{k}]$, where $\rho_{A}^{k}{=}\text{Tr}_{B}[\left|E_{k}\right\rangle\left\langle E_{k}\right|]$ is the reduced density matrix of subsystem $A$, obtained after tracing out subsystem $B$. For clarity, we denote the EE in the first case with equal subsystems $A$ and $B$ as $S$ and for the second case with non-equal subsystems $A$ and $B$ as $S_{\mathrm{mbl}}$.
In the ergodic phase, eigenstates are expected to resemble random pure states and exhibit volume-law EE consistent with Page entropy $S_{\mathrm{P}}{=}{\sum_{k=\mathcal{H}_{B}+1}^{\mathcal{H}_{A}\mathcal{H}_{B}}\ \frac{1}{k}{-}\frac{(\mathcal{H}_{A}-1)}{2\mathcal{H}_{B}}}$, wherein $\mathcal{H}_{A}$ ($\mathcal{H}_{B}$) is the dimension of the Hilbert space associated with the subsystem AA ($B$) [80, 81]. Studies indicate that highly excited eigenstates in the MBL phase obey an area law for bipartite entanglement, due to the emergent integrability [3, 21]. Therefore, averaged EE over eigenstates and disorder realizations $\overline{\langle S\rangle}$, results in $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}1$ in the ergodic phase, and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}0$ in the MBL phase.
@@
Figure 2: First row: the averaged EE as a function of disorder strength $W$ in (a) fully disordered chains, and in (b-d) disordered chains with thermal regions of different sizes $P$ (indicated in the insets). The system sizes $L$ used in numerical simulation are indicated in the insets. Second row: the averaged GR as a function of disorder strength $W$ for the same set of $P$ values. The results for EE and GR are obtained using 100020001000{-}2000 and 100002500010000{-}25000 realizations, respectively.
@@
The interplay between the MBL and ergodic phases is determined by two competing factors: (i) the disorder strength $W$; and (ii) the size of the thermal region $\mathcal{P}$, indicated by $P$ spins. While increasing $W$ tends to localize the system, increasing $P$ weakens the effect of disorder and tends to thermalize the system. In order to investigate these effects, we first fixed $P$ and vary $W$. By increasing $W$, the system goes through a phase transition from the ergodic to the localized phase. Figures 2(a)-(d) depict the behavior of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ as a function of the disorder strength $W$, for different system sizes $L{\in}[10,{\cdots},16]$, and various $P$’s. As Fig. 2(a) shows, the system is ergodic for small disorders and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}1$, indicating that the EE fulfills the volume-law entanglement, while increasing disorder $W$ leads to localization and $\overline{\langle S\rangle}/S_{\mathrm{P}}{\rightarrow}0$ [82]. The transition between these two phases happens at $W{=}W_{c}$. Thanks to scale invariant behavior at the phase transition point, one can utilize finite-size scaling analysis [83, 84, 85] to extract the critical value $W_{c}{\simeq}4.674\pm 0.308$ as the prediction for the transition point in the thermodynamic limit, for a system with P=0P{=}0. By embedding a thermal region, the critical disorder strength increases as can be noticed in Figs. 2(b)-(d) from the shift of the crossing points of the curves. Similar analysis can be performed for the GR which is shown in Figs. 2(e)-(h). Decreasing GR from $\overline{\langle r\rangle}{\simeq}0.5307$ to $\overline{\langle r\rangle}{\simeq}0.3863$ clearly demonstrates the transition from the ergodic to the MBL phase. Here, finite-size scaling analysis results in $W_{c}{\simeq}4.871\pm 0.243$ (for P=0P{=}0), which is close to the value identified by the EE. Repeating the above analysis, using either EE or GR, we extract $W_{c}$ for various values of $P$ and report the results in the Appendix (A). A shift in $W_{c}$ with increasing $P$ is noticeable in the case of GR as well.
Figure 3: (a) The averaged EE $\overline{\langle S\rangle}$, computed for half chain slicing as a function of P/LP{/}L. (b) the averaged EE $\overline{\langle S_{\mathrm{mbl}}\rangle}$, computed by separating the strongly disordered region on the left side from the rest of the chain as a function of P/LP{/}L. In both panels, the size of the system varies and W=9W{=}9. (c) The difference between the results presented in (a) and (b). (d) The averaged GR as a function of P/LP{/}L for W=9W{=}9. The results for EE and GR are obtained using 100020001000{-}2000 and 100002500010000{-}25000 realizations, respectively.
Alternatively, one can fix disorder strength $W$ to a large value, enough for localizing the system for P=0P{=}0, and then increase $P$ to thermalize the system gradually. In Fig. 3(a), we plot and $\overline{\langle S\rangle}/S_{\mathrm{P}}$ as a function of P/LP{/}L, keeping W=9W{=}9 fixed, for various system sizes. Several important features are observed. First, regardless of the system size, MBL can survive the presence of a small weak-disorder region $\mathcal{P}$, evidenced by the early plateau of the curves. Second, after the initial plateau when P/L>0.1P{/}L{>}0.1, the value of EE increases linearly by increasing $P/L$, such that the system eventually thermalizes to a volume law state. Third, for 0.1<P/L<0.30.1{<}P{/}L{<}0.3 the rate of change in the averaged EE diminishes by enlarging the system, ensuring the robustness of the MBL in the thermodynamic limit for the thermal region of size P>1P{>}1. Fourth, the rise of EE for P/L>0.3P{/}L{>}0.3, is sharpened as the system size increases, suggesting that in the thermodynamic limit, thermalization occurs abruptly. In the calculation of $\overline{\langle S\rangle}$, we put the cut in the middle of the chain, which is inside the region $\mathcal{P}$, and one might be concerned that it captures only the EE in the low-disordered region, caused by the thermalization in this area. To address this issue, we compute $\overline{\langle S_{\mathrm{mbl}}\rangle}$ by dividing the chain into non-equal subsystems, as illustrated on the right side of Fig. 1(c). This captures the EE behavior between the low-disordered area $\mathcal{P}$ and the strongly disordered part of the chain; see Fig. 3(b). Interestingly, $\overline{\langle S_{\mathrm{mbl}}\rangle}$ follows $\overline{\langle S\rangle}$ qualitatively. It shows the robustness of the MBL for P/L<0.1P{/}L{<}0.1, slower growth of the EE for 0.1<P/L<0.30.1{<}P{/}L{<}0.3, followed by sharper thermalization as the system size increases. In Fig. 3(c), we plot $\overline{\langle S\rangle}{-}\overline{\langle S_{\mathrm{mbl}}\rangle}$, which shows the contributions to EE from interactions between $P$ spins, due to the thermalization. As one can see, this portion of the EE shows the same behavior as $\overline{\langle S\rangle}$. This analysis ensures that, although entanglement inside the thermal region is involved, $\overline{\langle S\rangle}$ can perfectly reveal the behavior of the system in terms of P/LP{/}L and $W$.
@@
In Fig. 3(d), we depict $\overline{\langle r\rangle}$ versus P/LP{/}L in systems with different sizes, keeping the disorder strength fixed at W=9W{=}9. The results provide further qualitative confirmation for the findings from the EE. Initially, the GR increases very slowly indicating the resilience of the MBL phase against small weak-disorder regions in all considered system sizes. The growth of the GR increases significantly as P/LP{/}L increases, again indicating the instability of the MBL for reasonably large weak-disorder regions. Similar to the EE case, the growth rate sharpens as the system size increases. The above observations, for both GR and EE, lead us to the following conjecture: to trigger an avalanche and destabilize the MBL phase, the ratio $P/L$ should exceed a critical threshold, i.e. P/L>0.1P/L{>}0.1.
@@
To illustrate a complete picture of the effect of both disorder strength and the weak-disorder region, in Fig. 4(a) we plot $\overline{\langle S\rangle}/S_{\mathrm{P}}$ versus $W$ and P/LP{/}L, obtained in a system of size L=16L{=}16. The light yellow area represents the ergodic phase and the dark blue determines the MBL spot. As mentioned above, there are two ways to obtain the phase boundary (i) we can fix $P$ and vary $W$; or (ii) we can fix disorder strength $W$ and vary $P$. In the first approach, we find the intersection of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ curves for system of sizes $L$ and L+2L{+}2 when they are plotted as a function of $W$ for a given $P$, e.g. see Fig. 2(a) for P=0P{=}0. The intersection point $W_{c}^{*}(L_{av})$, with $L_{av}{=}L{+}1$ is taken as the finite-size transition point, at a given $P$, see Appendix (B) for more details. The results for $W_{c}^{*}(L_{av})$ are shown in Fig. 4(a) with solid circles. Different lines correspond to different sets of system sizes that are used to extract $W_{c}^{*}(L_{av})$. The smooth dependence of the EE on $L$ and $W$, especially at P=0P{=}0, guarantees that $W_{c}^{*}(L_{av})$ is a lower bound for the MBL transition point in the thermodynamic limit, namely $W_{c}$. As it is evident, by increasing $P$, the phase boundary initially remains straight and then bends after P/L>0.25P{/}L{>}0.25. This shows that the MBL phase is indeed stable against the presence of small thermal regions. On the other hand, one can also determine the phase boundary between the ergodic and the MBL phases by fixing $W$ and varying $P$. For every given $W$, we determine the transition point $(P{/}L_{av})_{c}^{*}$ by finding the intersection point of $\overline{\langle S\rangle}/S_{\mathrm{P}}$ curves for systems of size $L$ and L+2L{+}2, when they are plotted as a function of $P/L$ for a fixed value of $W$, e.g. see Fig. 3(a) for W=9W{=}9. The results are shown by stars in Fig. 4(a). Interestingly, the phase boundary between the MBL and the ergodic phase is found to be different depending on the two above approaches. This suggests that there exists an intermediate region between the ergodic and the MBL phase, which we cannot decisively label as MBL or ergodic. We call this area in the phase diagram the intermediate phase, which is also schematically shown in Fig. 1(b). As discussed for Figs. 3(a) and (d) (see also Figs. 11 and 12 in the Appendices for other $W$ values), this intermediate phase tends to vanish in the thermodynamic limit as the localization to ergodic transition accelerates by increasing the system size.
Figure 4: (a) averaged EE $\overline{\langle S\rangle}{/}S_{P}$ and (b) averaged GR $\overline{\langle r\rangle}$ as a function of $W$ and P/LP{/}L in a system of size L=16L{=}16. In both panels, the solid circles represent the estimated $W_{c}^{*}(L_{av})$ for fixed $P$’s. The other colorful markers represent the extracted $(P{/}L_{av})_{c}^{*}$ for fixed $W$’s. As graphically visualized in Fig. 1(b), three distinct regions are classified as $1.$ ergodic, $2.$ intermediate, and $3.$ localized.
@@
Having elucidated the EE behavior for inspecting the stability of MBL, we now focus on the GR. In Fig. 4(b) we report the obtained $\overline{\langle r\rangle}$ as a function of $W$ and P/LP{/}L for a system of size L=16L{=}16. By repeating the above analysis to determine the phase boundaries, one again finds different phase boundaries depending on whether $W$ is varied while $P/L$ is kept fixed or $P/L$ is varied while $W$ is fixed. Similar to the EE, the GR also shows relative stability for the MBL phase for small values of $P/L$, which is evident in the bending of the phase boundary as $P/L$ increases. Clearly, the GR analysis also leaves an intermediate phase between the ergodic and the MBL phases, which is not fully decisive. Therefore, both EE and GR predict an intermediate phase for finite systems which lies between the MBL and the ergodic phases.
IV Dynamical analysis